OpenStudy (anonymous):

What is the center and radius of the circle with the equation x2 + y2 + 8x + 4y – 61 = 0

OpenStudy (anonymous):

center is (-4,2) with radius 9

OpenStudy (anonymous):

@FlyinSolo_424 http://openstudy.com/study#/updates/4fc79d25e4b022f1e12f0873

OpenStudy (anonymous):

whats that link

OpenStudy (anonymous):

youll see

OpenStudy (anonymous):

its the awnser

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (mathstudent55):

x^2 + y^2 + 8x + 4y - 61 = 0 You need to complete the square for both x and y. (x^2 + 8x + ) + (y^2 + 4y + ) = 61 (x^2 + 8x + 16) + ( y^2 + 4y + 4) = 61 + 16 + 4 (x + 4)^2 + (y + 2) = 81 (x + 4)^2 + (y + 2)^2 = 9^2 Center: (-4, -2), radius = 9

OpenStudy (anonymous):

Basically get the 'x' terms & 'y' terms together comp,letetetete the square whaen everythings = to1 GL ;D